• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ole Hansen

    Log entry time 08:00:40 on June21,2000

    Entry number 44499

    keyword=Shift summary

    Started shift with production running at (-90,180) kinematics.

    1:40 MCC takes beam away for a few minutes for beam quality checks since Hall B is complaining.
    1:55 Beam returns. 94uA.
    2:02 Beam trips, MCC trying to improve things. OTR check.
    3:18 Beam dies, RF cavity problem.
    3:35 Beam is back.
    4:20 Flaky beam for a while (3-5 minutes), eventually stabilizes.
    4:40 More beam trips. Not very stable. MCC still sending 89uA.
    6:30 Current slowly dropping. Now at 86uA.
    7:00 Beam is more stable now. Trips are rare.
    7:05 ROC14 dies and takes runcontrol with it. Restarting CODA.
    7:11 Beam off. RF trip, qe checks.
    7:17 Beam back.
    7:25 The beam trips are back
    7:55 Stable beam now

    Mostly smooth running. Beam somewhat unstable for much of the night.


    Run Events Comments

    2354 1.3M 100uA current!
    2355 1.3M 50uA for a few minutes
    2356 1.3M
    2357 1.3M
    2358 1.3M
    2359 360k Beam off for OTR tests (Hall B is unhappy)
    2360 1.3M 94uA
    2361 1.3M
    2362 1.3M
    2363 1.3M
    2364 42k Short run, beam off
    2365 1.3M
    2366 1.3M
    2367 1.3M Frequent beam trips
    2368 1.3M
    2369 1.3M
    2370 1.3M
    2371 1.3M 86uA
    2372 1.3M
    2373 1.3M
    2374 370k ROC14 crash, short run
    2375 290k Beam off
    2376 1.3M
    2377 1.3M 84uA

    about 2.2C this shift